OpenStudy (anonymous):

Medal A chemistry experiment requires 14 ml of an acid solution. How many liters of acid should be ordered when 4 classes of 90 students each are going to perform the experiment? (The acid must be ordered by the whole liter.) ? L

OpenStudy (briensmarandache):

is that the whole question?

OpenStudy (jhannybean):

First change ml \(\to\) L

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

\[14~\text{mL} \cdot \frac{10^{-3}~\text{L}}{1~\text{mL}}\]

OpenStudy (jhannybean):

Then just multiply everything together, I believe.

OpenStudy (anonymous):

multiply what

OpenStudy (jhannybean):

Your conversion x # of classes x # of students.

OpenStudy (briensmarandache):

i mean one should do it, there is a 1000 ml in one liter, so one should handle four classes

OpenStudy (anonymous):

oh yeah

OpenStudy (princeharryyy):

as 4 classes=> 90*4*14=5040ml 1l = 1000 ml=> 40ml = 40/1000 = 0.4l total 5040 ml => 5l + 0.4l = 5.4l answer @harz360

OpenStudy (anonymous):

thank you soooo much :)

OpenStudy (princeharryyy):

welcome!

OpenStudy (princeharryyy):

@harz360